# Generative model for financial time series trained with MMD using a signature kernel

### v3

@@ 4.2 The Cell State and Hidden State of the LSTM
We emphasize that the conditioning described above is different from generating a conditional distribution such as in [37, Section 5] and [27, Section 3.4]. In both cases, the generator is trained to reproduce the conditional distribution of the future prices (or returns) given a specific past sequence. In [37], this is done by first assuming the true conditional expectation of the future price signature is a linear function of the past price signature and then using linear regression to estimate the parameters. Thereafter, the Monte Carlo estimate of the $l^{2}$ norm between the generator’s prediction given the past price serves as the loss function to train the generator. Using a linear estimator for the future path signature is no longer possible when we first lift the price path into a RKHS such as in our approach and in [27]. Therefore, the approach in [27] is to use a single historical sample of past and future sequence to estimate the score of multiple generated sequences conditioned on the same past sequence. In our approach, we estimate the score (MMD) by comparing a batch of generated sequences conditioned on a range of past sequences to the batch of future sequences resulting from the set of past sequences. We found that this approach combined with our choice of noise distribution described in Section 4.3 works well when generating long sequences (equal to about 1 calendar year of trading days) as we will show in Section 5.
@@ 4.3 Noise
For $\omega,\beta_{1},\ldots,\beta_{p}\in{\mathbb{R}}$ and $p\in{\mathbb{N}}$, the moving average model of order $p$ (MA$(p)$) is defined as
@@
Before performing the inverse Lambert W transformation, we perform a preprocessing step for numerical reasons explained in [35]. First, we annualise the log returns by dividing them by the time elapsed in years. Then we normalise them to have zero mean and unit variance to get the annualised and normalised returns, $r_{Z}$. The inverse Lambert W transformation is then applied by estimating the parameters $\delta,\mu_{R}$ and $\sigma_{R}$ using iterative generalised method of moments [^9] Although $r_{Z}$ has been normalised with zero mean and unit varaiance. The estimation process will aim to produce a kurtosis of 3 for the transformed variables which will cause the value of $\mu_{R}$ and $\sigma_{R}$ to deviate from zero. from [18, Section 4.2]. Using the estimated parameters, we obtain the "Gaussianised" returns, $r_{W}=W(r_{Z})$.
The parameters $\omega$ and $\beta_{1},\ldots,\beta_{p}$ of the MA model are then fitted using maximum likelihood estimation with $r_{W}$ as the target variable. When simulating noise for the generative model, we will sample a date from the historical data and label it $t_{0}$. The prior $p-1$ dates from the historical data will be labeled $t_{1-p},\ldots,t_{-1}$, where $p$ is the order of the MA model.
@@
Let ${\mathcal{H}_{z}}=(r_{W,t_{1-p}},\ldots,r_{W,t_{0}})$ be the historical data. Given parameters $\eta=\{\omega,\beta_{1},\ldots,\beta_{p}\}$ of the MA model and $d_{z}$ for the noise dimension, the noise at time $t_{i}$,
${\mathbb{R}}^{d_{z}}\ni z_{t_{i}}=(z_{1,{t_{i}}},\ldots,z_{d_{z},{t_{i}}})=:f_{\eta}({\mathcal{H}_{z}})\ \text{for}\ i\in\{1,\ldots,n\}$ (23)
is defined on the following operations:
@@
where $\epsilon_{t_{i}}$ is a standard normal random variable and $(z_{j,t_{1-p}},\ldots,z_{j,t_{0}})={\mathcal{H}_{z}}$ for $j\in\{1,\ldots,d_{z}\}$.
In other words, for the first $p$ time steps, the historical data is used in the generative process.
In summary, the goal of using a MA model, fitted in this manner, is to produce noise that is closer to the distribution of the historical returns than i.i.d. Gaussian noise. Based on the Heston process example, it has the potential of improving the performance of the generative model. The more difficult task of mapping from the noise to the empirical log returns will be done by the neural network.
@@ 4.4 Training Algorithm
With all the components in place, we can now describe the training algorithm for the generative model in Algorithm 1. Details of the implementation can be found in https://github.com/luchungi/Generative-Model-Signature-MMD.
Input: Initial generative model parameters $\Theta=\{\theta,\psi\}$, learning rate $\alpha$, MA model parameters $\eta$, Dataset ${\mathcal{D}}$, number of epochs $N$, batch size $B$, truncation level $m$, number of historical data points for the conditioning phase $k$, sample length $(n+1)$ and noise dimension $d_{z}$
@@ 4                   Sample a date from ${\mathcal{D}}$ without replacement and label it as $t_{0}$;
5                   Retrieve the historical data ${\mathcal{H}_{z}}=(r_{W,t_{1-p}},\ldots,r_{W,t_{0}})$ and $\mathcal{H}=(r_{{\mathcal{D}},t_{1}},\ldots,r_{{\mathcal{D}},t_{k-1}},\Delta t_{1},\ldots,\Delta t_{n})$ from ${\mathcal{D}}$;
6                   Sample a batch of noise $z_{t_{i}}=f_{\eta}({\mathcal{H}_{z}})$ for $i\in\{1,\ldots,n\}$ according to equation (23);
@@ 5.2 Noise Distribution
We test the conjecture that using noise which is closer to the distribution of the returns will improve the performance of the generative model with a toy example utilising the Heston stochastic volatility model introduced in [23]. We generate training data using the Heston model and fit two generative models. The first model uses i.i.d. standard Gaussian noise and the second model uses zero mean Gaussian noise with stochastic variance based on the same parameters as the simulation but independently generated. Details of the experiment parameters can be found in Appendix A.
@@ 5.3 Hyperparameters
We fit the MA model with maximum likelihood estimation using the python package arch [49]. We found empirically by a hyperparameter search that the MA$(20)$ model i.e. 20 lags produces good results.
@@
For the noise input, there is a choice for the dimension of the noise $d_{z}$. We investigated $d_{z}=1,2,4$ and found that $d_{z}=4$ produced the best results. The better performance when using a noise dimension of 4 could be due to a larger number of parameters in the model since a larger input size creates more parameters in the first layer of the neural network. The larger number of parameters could allow the model to better capture the complexity of the data. Therefore, to create a level playing field, we also experimented with quadrupling the hidden size to see if the performance gap between the noise dimensions would be reduced.
Table 2 shows the results of hypothesis testing with 100 generated samples, using the same permutation based two sample test as in Section 4.3 for different noise dimensions and hidden sizes. While enlarging the model size does improve the performance, there is still a significant performance gap between $d_{z}=4$ and other values. Enlarging the model with $d_{z}=4$ does not have any meaningful improvement in performance hence we choose the model with $d_{z}=4$ and hidden size of 64 as the final model.
@@
We demonstrate the impact on the $\widehat{\text{MMD}^{2}}$ by generating 100 batches of samples from the trained model. For each sample, we generate an alternative sample where we set all but the dominant noise dimension to zero. We then calculate the $\widehat{\text{MMD}^{2}}$ between the original S&P samples and the two generated samples. Table 3 shows the mean and standard deviation of the $\widehat{\text{MMD}^{2}}$ for the two types of samples. It is clear that the $\widehat{\text{MMD}^{2}}$ is significantly higher when all but the dominant noise dimension is set to zero despite the seemingly small differences in the generated log returns.
Sample Type $\widehat{\text{MMD}^{2}}$ mean $\widehat{\text{MMD}^{2}}$ std
Unaltered noise -0.472 3.249
Dominant dim 12.409 6.573
Table 3: Impact of noise on $\widehat{\text{MMD}^{2}}$. The $\widehat{\text{MMD}^{2}}$ for the dominant noise dimension is calculated using the noise sequence of the bottom left plot in Figure 5 which produces the log returns closest to output with the unaltered noise sequence.
@@ 5.5 Sensitivity Analysis
We studied sample sizes of 593,000 and 588,000, then rounded the extreme values to set the range of perturbation to $[-7.5,7.5]$ and $[-0.15,0.15]$ for the noise and returns respectively. Note that there were extreme outliers in the samples as 99.9% of the samples were within the range of $[-2.88,2.91]$ and $[-0.0543,0.0436]$ for the noise and returns respectively. Although we know that $\Delta t_{i}$ has an impact on the performance of the model from Section 5.4, the sensitivity is extremely low hence is omitted.
@@
Figure 7(a) shows the sensitivity of the output (log return) with respect to the noise and the previous return. For clarity, we only show the figure for the period from 18 Aug to 27 Oct 2008121212Omitted figures can be found at https://github.com/luchungi/Generative-Model-Signature-MMD/tree/main/figures. More importantly, it is apparent that the noise has a much larger impact on the output than the previous return. As it is difficult to see the relationship between the output and the previous return from Figure 7(a), we plot out the cross sections at fixed noise levels in Figure 7(b). Interestingly, the sensitivity to the previous return is larger at higher absolute noise levels with more positive returns resulting in larger absolute returns and vice versa.
@@ 5.6 Stylised Facts
As a comparison, we train a COT-GAN and Conditional Sig-WGAN (C-Sig-WGAN) generative models introduced in [59] and [37] respectively using the same data. We also use the same MA model to generate the noise input to the COT-GAN model, which has the same architecture including the number of noise dimensions. To ensure a fair comparison, we also use the same noise sequence for both models to generate the samples of the same length. The Conditional Sig-WGAN model is set up with the original optimised specification in [37] except with the signature truncation order increased to 5 which was the limit of the GPU memory available for the algorithm to run. We modify the code from the authors’ repository available at https://github.com/tianlinxu312/cot-gan for COT-GAN and https://github.com/SigCGANs/Conditional-Sig-Wasserstein-GANs for Conditional Sig-WGAN for our specific use case.
@@ 5.6.1 Moments of Returns
It is clear that the there can be significant variations in the statistics of the segments compared to the full sequence of S&P 500. Since the training of the generative model is based on the segments, it is not surprising and also desirable that the statistics of the generated sequences follow closer to the distribution of the segments.
@@
In addition, we also include the statistics of the moments for both our MMD based generator, the COT-GAN generator and the C-Sig-WGAN generator in Table 6. The table shows the mean and standard deviation of the distribution of the moments across 10,000 samples. In general, the models more or less capture the characteristics of the S&P 500. One deviation is in the samples from the C-Sig-WGAN which on average produces a slightly positive skew. The COT-GAN samples have a much wider distribution of the first moment compared to the other models and the S&P 500 segments. This is also evident when we look at the distribution of the end points of the generated segments.
@@
Figure 10 shows the density plot of the generated sequences’ end points with the initial points normalised to 1. Note that we truncated the x-axis for better visualisation as 35 of the COT-GAN samples had end points scattered between 2.5 and 5.6. The end points of the MMD generated sequences are more aligned with the S&P 500 segments than the COT-GAN generated sequences. The presence of more extreme returns explains the higher 2nd moment and lower 4th moment of the COT-GAN generated sequences due to a fatter middle. The extreme values of the COT-GAN generated sequences are overly pronounced and are not a realistic reflection of the training data. For the C-Sig-WGAN generated sequences, it is closer to the training data but the MMD generated sequences have a better overall fit.
@@
(b) COT-GAN
(c) C-Sig-WGAN
Figure 10: Density plot of the end points of the generated sequences. Initial values are normalised to start from 1.0 for all sequences. The x-axis is truncated for COT-GAN (due to more extreme vales). The S&P 500 histogram is constructed based on the training data construction method which are the same for MMD and COT-GAN but slightly different for C-Sig-WGAN which has shorter lengths.
@@ 5.6.2 Autocorrelation in Returns
In Figure 11(a), the statistics are consistent across the 3 types of sequences where it shows a lack of linear autocorrelation. However, returns are certainly not independent due to the phenomenon of volatility clustering. There are different manifestations of volatility clustering and one of them is the autocorrelation of the squared returns. Figure 11(b) shows the autocorrelation of the squared returns. While both the S&P 500 segments and full sequence exhibit positive and slow decaying autocorrelation with respect to an increasing time lag, the full sequence shows a more pronounced effect. The generated sequences are a closer match to the sampled sequences than to the full sequence as expected due to the training procedure.
(a) Autocorrelation of returns
(b) Autocorrelation of squared returns
Figure 11: Linear and non-linear autocorrelations of daily log returns of varying lags. Left and right figures show autocorrelation coefficient of the returns and squared returns, respectively. The blue and orange lines are the mean of the coefficient across the generated and sampled sequences, respectivel, while the shaded region shows the range of the mean absolute deviation. The red line is the coefficient of the full S&P 500 sequence.
Comparing the MMD generator with the COT-GAN and C-Sig-WGAN generators in Figure 12, we see that the MMD and COT-GAN generators produce similar autocorrelation of the squared returns in Figure 12(a). The C-Sig-WGAN generated sequences exhibit a much faster and smoother decay in the autocorrelation of the squared returns as shown in Figure 12(b). We omit the comparison for linear autocorrelation of the returns as the COT-GAN and C-Sig-WGAN generators produce similar results. [^14] See footnote 12.
(a) MMD vs COT-GAN for squared returns
(b) MMD vs C-Sig-WGAN for squared returns
Figure 12: Autocorrelation of squared returns for MMD, COT-GAN and C-Sig-WGAN. Lines and shaded regions are the same as in Figure 11.

### v4

@@ 4.2 The Cell State and Hidden State of the LSTM
We emphasize that the conditioning described above is different from generating a conditional distribution such as in [36, Section 5] and [27, Section 3.4]. In both cases, the generator is trained to reproduce the conditional distribution of the future prices (or returns) given a specific past sequence. In [36], this is done by first assuming the true conditional expectation of the future price signature is a linear function of the past price signature and then using linear regression to estimate the parameters. Thereafter, the Monte Carlo estimate of the $l^{2}$ norm between the generator’s prediction given the past price serves as the loss function to train the generator. Using a linear estimator for the future path signature is no longer possible when we first lift the price path into a RKHS such as in our approach and in [27]. Therefore, the approach in [27] is to use a single historical sample of past and future sequence to estimate the score of multiple generated sequences conditioned on the same past sequence. In our approach, we estimate the score (MMD) by comparing a batch of generated sequences conditioned on a range of past sequences to the batch of future sequences resulting from the set of past sequences. We found that this approach combined with our choice of noise distribution described in Section 4.3 works well when generating long sequences (equal to about 1 calendar year of trading days) as we will show in Section 5.
@@ 4.3 Noise
For $\omega,\beta_{1},\ldots,\beta_{p}\in{\mathbb{R}}$ and $p\in{\mathbb{N}}$, the moving average model of order $p$ (MA(p)(p)) is defined as
@@
Before performing the inverse Lambert W transformation, we perform a preprocessing step for numerical reasons explained in [35]. First, we annualise the log returns by dividing them by the time elapsed in years. Then we normalise them to have zero mean and unit variance to get the annualised and normalised returns, $r_{Z}$. The inverse Lambert W transformation is then applied by estimating the parameters $\delta,\mu_{R}$ and $\sigma_{R}$ using iterative generalised method of moments [^9] Although $r_{Z}$ has been normalised with zero mean and unit variance, the estimation process will aim to produce a kurtosis of 3 for the transformed variables which will cause the value of $\mu_{R}$ and $\sigma_{R}$ to deviate from zero. from [18, Section 4.2]. Using the estimated parameters, we obtain the "Gaussianised" returns, $r_{W}=W^{-1}(r_{Z})$.
The parameters $\omega$ and $\beta_{1},\ldots,\beta_{p}$ of the MA model are then fitted using maximum likelihood estimation with $r_{W}$ as the target variable. When simulating noise for the generative model, we will sample a date from the historical data and label it $t_{0}$. The prior p1p-1 dates from the historical data will be labeled $t_{1-p},\ldots,t_{-1}$, where $p$ is the order of the MA model.
@@
Let the noise at time $t_{i}$ be denoted by $z_{t_{i}}\in{\mathbb{R}}^{d_{z}}$ where $d_{z}\in{\mathbb{N}}$ is the dimension of the noise. Let ${\mathcal{H}_{p}}(i)=(z_{t_{i+1-p}},\ldots,z_{t_{i}})\in{\mathbb{R}}^{d_{z}p}$ be the realised noise sequence from time $t_{i+1-p}$ to $t_{i}$ for $i\in\{1,\ldots,n\}$ and $p\in{\mathbb{N}}$ be the order of the MA model from Definition 12. Specifically, if $z_{j,t_{i}}$ is the $j$-th component of $z_{t_{i}}$, then let $z_{j,t_{i}}=r_{W,t_{i}}$ for $i\in\{1-p,\ldots,0\}$, $j\in\{1,\ldots,d_{z}\}$ where $r_{W,t_{i}}\in{\mathbb{R}}$ is the Gaussianised return at time $t_{i}$. Given parameters $\eta=\{\omega,\beta_{1},\ldots,\beta_{p}\}$ of the MA model from Definition 12, we define the noise at time $t_{i}$ via the map $f_{\eta}$ given by
${\mathbb{R}}^{d_{z}p}\ni(z_{t_{i-p}},\ldots,z_{t_{i-1}})={\mathcal{H}_{p}}(i-1)\mapsto f_{\eta}({\mathcal{H}_{p}}(i-1)):=z_{t_{i}}\ \text{for}\ i\in\{1,\ldots,n\}$ (23)
which is explicitly defined through the following operations:
@@
where $\epsilon_{t_{i}}$ is a standard normal random variable.
In other words, for the first $p$ time steps, the historical data is used in the generative process, thereafter, the noise is generated using the MA model based on the previous $p$ noise values. In our experiments reported in Section 5, we use $p=20$.
In summary, the goal of using a MA model, fitted in this manner, is to produce noise that is closer to the distribution of the historical returns than i.i.d. Gaussian noise. Note that we do not reverse the transformations, which include inverse Lambert W, annualisation and normalisation, to the sampled noise $z_{j,t_{i}}$ as we found that this did not work empirically. We speculate that this is due to the reverse process producing returns with a much more extreme fat tail than observed in the original data. Therefore, we leave the more difficult task of mapping from the noise to the empirical log returns to be done by the neural network.
@@ 4.4 Training Algorithm
With all the components in place, we can now describe the training algorithm for the generative model in Algorithm 1. Details of the implementation can be found in https://github.com/Anonymous1701/mmd-repo.
Input: Initial generative model parameters $\Theta=\{\theta,\psi\}$, learning rate $\alpha$, MA model parameters $\eta$ and $p$, Dataset $\mathcal{D}$}, number of epochs NN, batch size $B$, truncation level $m$, number of historical data points for the conditioning phase $k$, sample length $(n+1)$ and noise dimension $d_{z}$
@@ 4         Sample a date from $\mathcal{D}$} without replacement and label it as $t_{0}$;
5         Retrieve the historical data ${\mathcal{H}_{p}}(0)=(r_{W,t_{1-p}},\ldots,r_{W,t_{0}})$ and $\mathcal{H}=(r_{{\mathcal{D}},t_{1}},\ldots,r_{{\mathcal{D}},t_{k-1}},\Delta t_{1},\ldots,\Delta t_{n})$ from $\mathcal{D}$};
6         Sample a batch of noise $z_{t_{i}}=f_{\eta}({\mathcal{H}_{p}}(i-1))$ for $i\in\{1,\ldots,n\}$ according to equation (23);
@@ 5.2 Noise Distribution
We test the conjecture that using noise which is closer to the distribution of the returns will improve the performance of the generative model with a toy example utilising the Heston stochastic volatility model introduced in [23]. We generate training data using the Heston model and fit two generative models. The first model uses i.i.d. standard Gaussian noise and the second model uses zero mean Gaussian noise with stochastic variance [^10] The reader may wonder why we did not use the noise process described in Section 4.3. The main goal is to provide some empirical evidence that if the noise input for the generative model is closer to the actual distribution of the paths, it should produce a better result. In this particular case, we chose the noise distribution to follow the CIR process for the variance of the price process since it is the actual ”noise” process of the Heston model. based on the same parameters as the simulation but independently generated. Details of the experiment parameters can be found in Appendix A.
@@ 5.3 Hyperparameters
We fit the MA model with maximum likelihood estimation using the python package arch [48]. We found empirically by a hyperparameter search that the MA(20)(20) model i.e. 20 lags produces good results.
@@
For the noise input, there is a choice for the dimension of the noise $d_{z}$. We investigated $d_{z}=1,2,4,6,8$ and found that $d_{z}=4$ appears to be a minimum to achieve a good result in terms of convergence in the $\widehat{\text{MMD}^{2}}$. The better performance when using a noise dimension of 4 could be due to a larger number of parameters in the model since a larger input size creates more parameters in the first layer of the neural network. The larger number of parameters could allow the model to better capture the complexity of the data. Therefore, to create a level playing field, we also experimented with quadrupling the hidden size for $d_{z}=1,2,4$ to see if the performance gap between the noise dimensions would be reduced.
Table 2 shows the results of hypothesis testing with 100 generated samples, using the same permutation based two sample test as in Section 5.2 for different noise dimensions and hidden sizes. While enlarging the model size does improve the performance, there is still a significant performance gap between $d_{z}=4$ and the smaller values. Enlarging the model with $d_{z}=4$ or increasing the number of noise dimensions does not have any meaningful improvement in performance hence we choose the model with $d_{z}=4$ and hidden size of 64 as the final model.
@@
We demonstrate the impact on the $\widehat{\text{MMD}^{2}}$ by generating 100 samples from the trained model. For each sample, we generate an alternative sample where we set all but the dominant noise dimension to zero. We perform the same permutation based two sample tests as in Section 5.2 to compare the two sets of generated sequences with the original S&P 500 sequences. It is clear from Table 3 that despite the seemingly small differences in the generated log returns when all but the dominant noise dimension is set to zero, it has a significant impact on the $\widehat{\text{MMD}^{2}}$ measured against the S&P 500 samples.
Sample Type $\widehat{\text{MMD}^{2}}$ p-value
Unaltered noise -0.578 0.628
Dominant dim 12.239 0.0
Table 3: Impact of noise on the $\widehat{\text{MMD}^{2}}$ with 10,000 permutations. The unaltered noise is the noise used to train the model while the dominant dim is the noise with all but the dominant noise dimension set to zero.
@@ 5.5 Sensitivity Analysis
We studied sample sizes of 593,000 and 588,000, then rounded the extreme values to set the range of perturbation to [7.5,7.5][-7.5,7.5] and [0.15,0.15][-0.15,0.15] for the noise and returns respectively. Note that there were extreme outliers in the samples as 99.9% of the samples were within the range of [2.88,2.91][-2.88,2.91] and [0.0543,0.0436][-0.0543,0.0436] for the noise and returns respectively. Although we know that $\Delta t_{i}$ has an impact on the performance of the model from Section 5.4, the sensitivity is extremely low hence is omitted.
@@
Figure 7(a) shows the sensitivity of the output (log return) with respect to the noise and the previous return. For clarity, we only show the figure for the period from 18 Aug to 27 Oct 2008131313Omitted figures can be found at https://github.com/Anonymous1701/mmd-repo/tree/main/figures. More importantly, it is apparent that the noise has a much larger impact on the output than the previous return. As it is difficult to see the relationship between the output and the previous return from Figure 7(a), we plot out the cross sections at fixed noise levels in Figure 7(b). Interestingly, the sensitivity to the previous return is larger at higher absolute noise levels with more positive returns resulting in larger absolute returns and vice versa.
@@ 5.6 Stylised Facts
As a comparison, we train a signature kernel based neural SDE (NSDE), COT-GAN and Conditional Sig-WGAN (C-Sig-WGAN) generative models introduced in [27], [58], and [36], respectively using the same data. We also use the same MA model to generate the noise input to the COT-GAN model, which has the same architecture including the number of noise dimensions. To ensure a fair comparison, we also use the same noise sequence for both our model and the COT-GAN model to generate the samples of the same length. The Conditional Sig-WGAN model is set up with the original optimised specification in [36] except with the signature truncation order increased to 5 which was the limit of the GPU memory available for the algorithm to run. For the NSDE model, we use the same hyperparameters as in [27] for the unconditional setting except for the data related hyperparameters which are adjusted to match our use case. We also tried varying the $\sigma$ parameter of the RBF kernel in the NSDE model but found that it did not have a significant impact on the results hence we stuck to the original value of 1.0. We could not use the conditional setting of the NSDE model as the hyperparameters of the orignal model was unable to produce any meaningful results. We modify the code from the authors’ repository available at https://github.com/issaz/sigker-nsdes for NSDE, https://github.com/tianlinxu312/cot-gan for COT-GAN and https://github.com/SigCGANs/Conditional-Sig-Wasserstein-GANs for Conditional Sig-WGAN for our specific use case.
@@ 5.6.1 Moments of Returns
It is clear that there can be significant variations in the statistics of the segments compared to the full sequence of S&P 500. Since the training of the generative model is based on the segments, it is not surprising and also desirable that the statistics of the generated sequences follow closer to the distribution of the segments.
@@
In addition, we also include the statistics of the moments for both our MMD based generator, the NSDE generator, the COT-GAN generator and the C-Sig-WGAN generator in Table 6. The table shows the mean and standard deviation of the distribution of the moments across 10,000 samples. In general, the models more or less capture the characteristics of the S&P 500. One deviation is in the samples from the NSDE and C-Sig-WGAN models which on average produces a slightly positive skew. The NSDE model also does not produce fatter tails which is evident in the kurtosis of the distribution. The COT-GAN samples have a much wider distribution of the first moment compared to the other models and the S&P 500 segments. This is also evident when we look at the distribution of the end points of the generated segments.
@@
Figure 10 shows the density plot of the generated sequences’ end points with the initial points normalised to 1. Note that we truncated the x-axis for better visualisation as 35 of the COT-GAN samples had end points scattered between 2.5 and 5.6. The end points of the MMD generated sequences are more aligned with the S&P 500 segments than the COT-GAN generated sequences. The presence of more extreme returns explains the higher 2nd moment and lower 4th moment of the COT-GAN generated sequences due to a fatter middle. The extreme values of the COT-GAN generated sequences are overly pronounced and are not a realistic reflection of the training data. For the NSDE and C-Sig-WGAN generated sequences, it is closer to the training data but the MMD generated sequences have a better overall fit.
@@
(b) NSDE
(c) COT-GAN
(d) C-Sig-WGAN
Figure 10: Density plot of the end points of the generated sequences. Initial values are normalised to start from 1.0 for all sequences. The x-axis is truncated for COT-GAN (due to more extreme vales). The S&P 500 histogram is constructed based on the training data construction method which are the same for MMD, NSDE and COT-GAN but slightly different for C-Sig-WGAN which has shorter lengths.
@@ 5.6.2 Autocorrelation in Returns
In Figure 11, the statistics are consistent across the 3 types of sequences where it shows a lack of linear autocorrelation. We omit the comparison for linear autocorrelation of the returns as the NSDE, COT-GAN and C-Sig-WGAN generators produce similar results. [^15] See footnote 13.
Figure 11: Linear autocorrelations of log returns. The blue and orange lines are the mean of the coefficient across generated sequences, while the shaded region covers the mean absolute deviation. The red line is the coefficient of the full S&P 500 sequence.
However, returns are certainly not independent due to the phenomenon of volatility clustering. There are different manifestations of volatility clustering, and one of them is the autocorrelation of the squared returns. Figure 12 shows the autocorrelation of the squared returns for all generated sequences. While both the S&P 500 segments and full sequence exhibit positive and slow decaying autocorrelation with respect to an increasing time lag, the full sequence shows a more pronounced effect. The MMD and COT-GAN generated sequences are a closer match to the S&P 500 segments. The C-Sig-WGAN generated sequences exhibit a much faster and smoother decay in the autocorrelation of the squared returns. The NSDE generated sequences do not capture the autocorrelation of the squared returns well, with an almost flat autocorrelation curve.
(a) MMD
(b) NSDE
(c) COT-GAN
(d) C-Sig-WGAN
Figure 12: Autocorrelation of squared returns for all generated sequences. Lines and shaded regions are the same as in Figure 11.